OpenStudy (junodan):

can someone help me find the equation of the AOS of the graph of > y + 3x – 6 = –3(x – 2)2 + 4?

OpenStudy (junodan):

@mayankdevnani can you help me with this one as well?

OpenStudy (mayankdevnani):

What do you think ?

OpenStudy (junodan):

1/6 was my first guess

OpenStudy (junodan):

-1/6

OpenStudy (mayankdevnani):

woo !! that was wrong

OpenStudy (mayankdevnani):

to find it, use the formula :- x=-b/2a

OpenStudy (mayankdevnani):

this vertical line is our axis of symmetry

OpenStudy (junodan):

i used it but i think i messed up i got -1/2 and then i got 3/2

OpenStudy (junodan):

@mayankdevnani How bad did i mess these up lol

OpenStudy (mayankdevnani):

can you please simplfy your equation?

OpenStudy (junodan):

what do you mean?

OpenStudy (mayankdevnani):

y + 3x – 6 = –3(x – 2)2 + 4 simplify it to standard quadratic equation

OpenStudy (junodan):

oh ok hang on

OpenStudy (junodan):

im getting it

OpenStudy (junodan):

its ok ill just guess on this one mt teacher will explain it to me

OpenStudy (mayankdevnani):

y + 3x – 6 = –3(x – 2)2 + 4 y+3x-6=-3(x2+4-4x)+4 y+3x-6=-3x2-12+12x+4

OpenStudy (mayankdevnani):

simplify it